# TopoGEN: topology-driven microstructure generation for in silico modeling of fiber network mechanics

### v1

@@ Abstract
The fields of mechanobiology and biomechanics are expanding our understanding of the complex behavior of soft biological tissues across multiple scales. Given the intricate connection between tissue microstructure and its macroscale mechanical behavior, unraveling this mechanistic relationship remains an ongoing challenge. Reconstituted fiber networks serve as valuable in vitro models to simplify the intricacy of in vivo systems for targeted investigations. Concurrently, advances in imaging enable microstructure visualization and, through generative pipelines, modeling as discrete element networks. These mesoscale (µm) models provide insights into macroscale (mm) tissue behavior. However, a systematic study of how microstructural variations influence nonlinear tissue mechanics is still lacking. In this work, we develop an innovative framework to generate topologically-driven discrete fiber networks. Leveraging these networks, we generate models of interconnected load-bearing fiber components that exhibit softening under compression and are bending-resistant. By virtually replicating microstructural features of reconstituted collagen networks, such as fiber volume fractions and cross-link concentration, we evaluate the robustness of our simulations. Analyzing the nonlinear elastic behavior at varying polymerization temperatures, we find consistency between the in silico results and in vitro data from the literature. We extend our investigation beyond empirically measurable factors to explore microstructural effects at the single fiber level (i.e., fibril morphology and stiffness) that are challenging to investigate experimentally. TopoGEN allows us to mechanistically explore localized microstructural phenomena and relate microstructural changes to the bulk mechanical response of soft biological materials, hence providing an indispensable tool to further advance the fields of tissue biomechanics and engineering.
@@
[label1]organization=Department of BioMechanical Engineering, Faculty of Mechanical Engineering,
Delft University of Technology, country=the Netherlands
@@ 1 Introduction
Microstructure drives function. In soft biological materials, the microstructure comprises a complex fibrous network known as the extracellular matrix. This intricate three-dimensional meshwork of fibers provides mechanical stability, elasticity, and strength and plays a crucial role in governing cellular processes (Dean et al., 2023). Cells interact with the extracellular matrix through mechanotransduction (Jansen et al., 2015), a process by which they sense environmental changes and convert mechanical impulses into biological responses (Fereidoonnezhad et al., 2017; Peirlinck et al., 2019). The discipline of mechanobiology revolves around this dynamic feedback loop (Loerakker and Humphrey, 2022), which is regulated by highly localized micromechanical factors. A key challenge, shared across many engineering applications (Geers et al., 2010), is understanding the relationship between microstructural features (e.g., individual constituents and network properties) and macroscopic functions (e.g., mechanical behavior) of these networks. In the specific case of soft biological tissues, such as collagen network within the extracellular matrix, the primary challenge lies in the geometrical complexity and weakly connected microstructure that disrupts deformation affinity and gives rise to localized deformations (Mirzaali et al., 2020) highlighted in experimental studies (Cavinato et al., 2020). Therefore, in biological tissues, the relative spatial position of two neighboring particles after deformation cannot be fully described solely by their relative material position before deformation, and non-affinity must be considered. Addressing this challenge requires in-depth investigation of details associated with the individual constituents of the networks, such as fiber kinematics (Mahutga et al., 2023). Early computational models of soft biological tissues were predominantly continuum-based, drawing inspiration from histological data (Fung, 1967; Holzapfel et al., 2000; Gasser et al., 2005). These models rely on intuitive understanding, data availability, and a priori physical assumptions. Thus, they are descriptive tools of material mechanics but fall short in predicting complex loading scenarios or new materials behavior. Most notably, they fail to offer mechanistic insights into material behavior based on the structural information of their building blocks. Advancements in imaging techniques have led to more detailed representations of the extracellular space through network models. Such network models explicitly depict interconnected fibers as discrete networks (Lindström et al., 2010, 2013; Nan et al., 2018; Ban et al., 2019; Merson and Picu, 2020; Dalbosco et al., 2021; Eichinger et al., 2021a; Leng et al., 2021; Nikpasand et al., 2021; Filla et al., 2023; Kakaletsis et al., 2023; Mahutga et al., 2023; Wahlsten et al., 2023), establishing the connection between fiber mechanics and tissue behavior. On the experimental side, tissue engineers have developed quantitative techniques to precisely control the microstructure of reconstituted gels (e.g., by adjusting temperature or concentration) and assess their influence on mechanical response (Jansen et al., 2018). These analyses show that even small environmental changes can significantly impact microstructure. The bulk mechanical response results from multiple coexisting factors. Potential microstructural factors include fiber interconnectivity (i.e., the topology), fiber concentration in the sample volume, and their morphological properties. However, the interrelated and multiscale nature of these factors makes it challenging to disentangle the influence of individual variables on the overall mechanical behavior of the sample. Computational models that adopt discrete fiber networks offer a highly valuable complementary approach to experiments, enabling a systematic examination of how specific microstructural features contribute to macroscopic properties. Most of the current in silico discrete fiber-based approaches are concentration-driven rather than topology-driven, matching fiber volume fractions from experiments. However, this approach overlooks the critical role of network topology and other microstructural parameters, such as the morphological characteristics of individual fiber constituents, in shaping mechanical behavior. Understanding the localized deformations and varied mechanical properties of these reconstituted biological networks is crucial for uncovering how mechanical cues regulate cellular behavior, influence tissue development, and contribute to disease progression (van Putten et al., 2016; Herum et al., 2017; Ayad et al., 2019; Buskermolen et al., 2020; Eichinger et al., 2021b; Espina et al., 2021; Kim et al., 2021; Atcha et al., 2023; Kumar et al., 2024; McEvoy and McNamara, 2024). Moreover, investigating how the microstructure, in terms of fiber and network properties, relates to the measured properties enables material behavior prediction and provides a solid foundation for designing engineered tissues with tailored macroscopic mechanics (Burla et al., 2019b, 2020).
In this study, we propose TopoGEN, a framework that integrates three-dimensional fiber network architectures with nonlinear finite element analysis to systematically explore the intricate governing structure-function relationships in soft tissues. Our framework features a generative pipeline to create topologically rich and microstructurally varying discrete fiber networks. We explore the effect of changing the microstructure on the overarching macroscale behavior and compare our results to experimentally studied reconstituted collagen networks, where similar topology variations are induced through temperature effects. To extend our analysis beyond experimentally measurable mesoscale parameters, such as concentration, we investigate complex microstructural effects, including the morphology and mechanics of individual fiber constituents, on the bulk mechanical response of soft biological materials.
@@ 2.1 Fiber network generation
We model biological networks as structures of connected fibers. To generate these networks, we partition a three-dimensional cubic domain into randomly seeded Voronoi polyhedrons and represent the fibers and cross-links as the tesselation and vertices, respectively. We then optimize this Voronoi-based fiber network using simulated annealing optimization technique (Yeong and Torquato, 1998; Lindström et al., 2010, 2013; Eichinger et al., 2021a) to match key topological features, i.e., average connectivity (valency) and fiber length distribution, to those of experimentally measured collagen fiber networks (Jansen et al., 2018). We set the to-be-minimized annealing system energy as the distance of the initial microstructural configuration from the target microstructural configuration achieved by gradually cooling the network and allowing it to transition to new neighboring microstructures. These neighboring states are generated by applying small random perturbations to the configuration at each iteration until the current and target distributions match. To address the challenge of directly comparing continuous length distributions to the log-normal target (Lindström et al., 2010), we implement a binning algorithm that partitions fiber lengths into disjoint intervals ($b$). Therefore, we assign each length observation ($x$) to a specific interval and compute a discretized cost function expressing the difference between the binned distribution and the target. To this aim, we implement the Kullback–Leibler (KL) divergence (Kullback and Leibler, 1951), also known as relative entropy ($\mathcal{D}_{KL}$), which is a non-symmetric measure quantifying how one probability distribution ($p(x)$) diverges from a second reference probability ($q(x)$) distribution:
@@
To transform our initial Voronoi-generated random fiber network into neighboring states at each simulated annealing iteration, we introduce two types of moves illustrated in Figure 1: dilutive transformations, which randomly remove fibers within the network and concentration-preserving transformations, which randomly move nodes within the cubic domain.
Figure 1: Simulated annealing workflow. Starting from an initial random configuration, the algorithm iteratively minimizes the system’s energy by removing fibers (move 1) or moving nodes (move 2) until the target distribution of valency and fiber length is achieved.
Here, we define cross-links as nodes with fourfold valency and branches as nodes with threefold valency. Starting from the fourfold valency typical of Voronoi networks, dilution allows branch formations until the target valency is reached. The Poisson length distribution used to generate the initial network results in a number of short fibers, which are inconsistent with experimentally measured target distribution. To address this, we assign higher removal probabilities to shorter edges during the dilutive transformations and iteratively adjust node position to match the log-normal distribution during the concentration-preserving transformations.
@@ 2.2 Mechanical equilibrium at the mesoscale
Modeling soft tissues at the level of individual fibers (microscale) or the level of discrete fiber networks (mesoscale) is computationally impractical for large volumes undergoing complex loading and boundary conditions. Instead of extending our discrete fiber network model to the entire tissue scale (macroscale), we focus on a representative volume element (RVE) large enough to statistically capture the mechanical properties of the microstructure yet small enough to reduce the computational cost. Within this mesoscale domain, we define a continuum-based constitutive behavior for each individual fiber in the network and apply macroscopic deformation gradients (i.e., tissue loading conditions) as boundary conditions. Solving this boundary value problem with the finite element method yields the macroscale’s local (i.e., for every finite element, at each integration point, within every time step, at each Newton iteration) constitutive law (Geers et al., 2010; Peirlinck et al., 2024). In this study, we use the commercial finite element analysis software Abaqus Dassault Systèmes Simulia Corp. (2024) to retrieve the numerical solution of the RVEs boundary value problem. The transition from the RVE to the full tissue scale is only admissible if the Hill-Mandel condition (Hill, 1963) is satisfied, ensuring that the potential energy remains consistent when passing from the meso- to the macroscale. This condition is ideally met with infinitely large RVEs; however, with finite-sized RVEs, the consistency requires careful handling of the boundary conditions and the RVE size. This can be achieved through three approaches: (1) enforcing zero microfluctuations throughout the RVE, i.e. coupling local RVE deformation affinely to the macroscale deformation gradient; (2) setting zero microfluctuations only at the boundary via uniform displacement boundary conditions while allowing internal nodes to deform non-affinely (i.e., with deviations from a direct one-to-one correspondence to the macroscale deformation gradient); or (3) imposing periodicity of the microfluctuation field at the boundary via periodic boundary conditions (PBCs) and allowing internal nodes to deform non-affinely. In this work, we adopt PBCs since they are known to estimate the overall material properties better than the other alternatives mentioned (Miehe and Koch, 2002; van der Sluis et al., 2000; Terada et al., 2000; Kanit et al., 2006; Perić et al., 2010) and lie closer to the effective properties compared to the upper and lower bounds defined by uniform displacement and traction boundary conditions. Another advantage of adopting PBCs is their ability to capture the effect of non-affinity on the RVE response by allowing internal nodes to move freely within our loosely connected networks. Given that biological networks are - in essence - not periodic structures, incorporating these PBCs is not a trivial task. Figure 2 highlights how we enforce periodicity at the boundary nodes while maintaining the internal microfluctuations. First, we randomly seed our RVE domain with N seeding points and generate 26 periodic replicas of this domain by offsetting the seeding points around the central domain (Figure 2, step 1). Next, we apply Voronoi tessellation using 27 X N points as seeding particles and extract the interconnected fibers from the skeletonized diagrams (Figure 2, step 2). We crop the larger network to the original cubic domain and assemble a periodic edge connectivity map between each boundary node and its periodic counterpart on the opposing edge (Figure 2, step 3). Finally, we implement PBCs by constraining the relative displacements of all node pairs, as defined in our assembled periodic boundary node connectivity table, to individual dummy nodes. These dummy nodes, unattached to any fiber in the model, serve as reference points used in the nodal equations that impose the periodicity of the nodally paired loading conditions (Wu et al., 2014). In particular, at each individual dummy node, we enforce an affine macroscopic deformation gradient corresponding to the macroscopic RVE face deformation profile under study (e.g., uniaxial extension, biaxial extension, or shear, as illustrated in Figure 2, step 4). For each of these macroscopic deformation profiles, we introduce a linear interpolation map to constrain the boundary nodes to remain planar during shear or uni/biaxial tests.
@@
Focusing purely on elastic effects in this work, we assume fiber connections to be permanent, assuming no cross-link or branch breakage. We consider that the persistence length of fibers, i.e., the characteristic length along the polymer chain over which its direction remains correlated before it starts to bend, is typically larger than the contour length. Therefore, the fibers are bending-stiff, and their thermal fluctuations can be neglected (Burla et al., 2019a). While the Euler-Bernoulli beam formulation could have been considered (Shahsavari and Picu, 2012), neglecting shear stiffness is mostly suited for fibers that are highly slender, i.e., with a cross-section-to-length ratio of less than 1:15. However, experiments in biological networks (e.g., collagen and fibrin) reveal a log-normal distribution of fiber lengths, which results in short fibers that violate this requirement (Lindström et al., 2010, 2013). Therefore, we model individual fiber mechanics using second-order Timoshenko shear-flexible beams, which are particularly suited for modeling biological networks (Shahsavari and Picu, 2013). We assume our fibers behave linearly elastic in both extension and compression, where we take into account fiber softening under compression by setting the compressive Young’s modulus to be a tenth of the tensile Young’s modulus (Notbohm et al., 2015). To incorporate this bilinear tension-compression asymmetry, we developed a custom user-defined field USDFLD subroutine that defines the fiber stiffness at each integration point based on the axial strain of the loaded beam. Upon solving the mechanical equilibrium of our topologically optimized RVEs under various periodic loading conditions, we obtain the reaction forces on our PBC-constrained faces. We use these reaction forces to compute a representative stress tensor across the entire discrete fiber network by calculating the first Piola–Kirchhoff stress tensor ($\mathbf{P}$):
@@
where $A_{0}$ denotes the initial surface of our cubic RVE, ${N_{i}}$ are the dummy nodes, and $\mathbf{f}_{i}$ represents the resulting external forces acting on these nodes.

### v2

@@ Abstract
The fields of mechanobiology and biomechanics are expanding our understanding of the complex behavior of soft biological tissues across multiple scales. Given the intricate connection between tissue microstructure and its macroscale mechanical behavior, unraveling this mechanistic relationship remains an ongoing challenge. Reconstituted fiber networks serve as valuable in vitro models to simplify the intricacy of in vivo systems for targeted investigations. Concurrently, advances in imaging enable microstructure visualization and, through generative pipelines, modeling as discrete element networks. These mesoscale (µm) models provide insights into macroscale (mm) tissue behavior. However, there is still no clear way to systematically incorporate detailed experimentally observed microstructural changes into in silico models of biological networks. In this work, we develop a novel framework to generate topologically-driven discrete fiber networks using high-resolution images that account for how environmental changes during polymerization influence the resulting structure. Leveraging these networks, we generate models of interconnected load-bearing fiber components that exhibit softening under compression and are bending-resistant. The generative topology framework enables control over network-level features, such as fiber volume fraction and cross-link density, along with fiber-level properties, like length distribution, to simulate changes driven by different polymerization conditions. We validate the robustness of our simulations against experimental data in a collagen-specific study case where we examine nonlinear elastic responses of collagen networks across varying conditions. TopoGEN provides a versatile tool for tissue biomechanics and engineering, helping to bridge microstructural insights and bulk mechanical behavior by linking image-derived microstructural topological organization to soft tissue mechanics.
@@
[label1]organization=Department of BioMechanical Engineering, Faculty of Mechanical Engineering, Delft University of Technology, country=the Netherlands
@@ 1 Introduction
Microstructure drives function. In soft biological materials, the microstructure comprises a complex fibrous network known as the extracellular matrix. This intricate three-dimensional meshwork of fibers provides mechanical stability, elasticity, and strength and plays a crucial role in governing cellular processes [20]. Cells interact with the extracellular matrix through mechanotransduction [42], a process by which they sense environmental changes and convert mechanical impulses into biological responses [27, 71]. The discipline of mechanobiology revolves around this dynamic feedback loop [54], which is regulated by highly localized micromechanical factors. A key challenge, shared across many engineering applications [32], is understanding the relationship between microstructural features (e.g., individual constituents and network properties) and macroscopic functions (e.g., mechanical behavior) of these networks. In the specific case of soft biological tissues, the primary challenge lies in the geometrical complexity and weakly connected microstructure that disrupts deformation affinity and gives rise to localized deformations [62] highlighted in experimental studies [13]. Therefore, in biological tissues, the relative spatial position of two neighboring particles after deformation cannot be fully described solely by their relative material position before deformation, and non-affinity must be considered. Addressing this challenge requires an in-depth investigation of details associated with the individual constituents of the networks, such as fiber kinematics [55]. Early computational models of soft biological tissues were predominantly continuum-based, drawing inspiration from histological data [29, 36, 31]. These models rely on intuitive understanding, data availability, and a priori made physical assumptions. Thus, they serve as descriptive tools of material mechanics but fall short in predicting complex loading scenarios or new material behavior [88, 16, 74, 90, 55]. Most notably, they fail to offer mechanistic insights into material behavior based on the structural information of their building blocks as existing continuum models cannot capture the non-affine fiber kinematics, resulting in predictions that are too stiff compared to experiments [74]. Advancements in imaging techniques have led to more detailed representations of the extracellular space through network models. Such network models explicitly depict interconnected fibers as discrete networks [53, 76, 52, 64, 5, 59, 17, 24, 49, 65, 28, 44, 55, 93], establishing the connection between fiber mechanics and tissue behavior. On the experimental side, tissue engineers have developed quantitative techniques to precisely control the microstructure of reconstituted gels by adjusting temperature or pH [81, 79, 80, 38, 94, 82, 63, 57, 67, 33, 90, 91] and assess their influence on mechanical response [43]. These analyses show that even small environmental changes can significantly impact microstructure. The bulk mechanical response results from multiple coexisting factors. Potential microstructural factors include fiber interconnectivity (i.e., the topology), fiber concentration in the sample volume, and their morphological and mechanical properties. However, the interrelated and multiscale nature of these factors makes it challenging to disentangle the influence of individual variables on the overall mechanical behavior of the sample. Computational models that adopt discrete fiber networks offer a highly valuable complementary approach to experiments, enabling a systematic examination of how specific microstructural features contribute to macroscopic properties. Most of the current in silico discrete fiber-based approaches are concentration-driven rather than topology-driven, matching fiber volume fractions from experiments. However, this approach overlooks the critical role of network topology and other microstructural parameters in shaping mechanical behavior. Understanding the localized deformations and varied mechanical properties of these reconstituted biological networks is crucial for uncovering how mechanical cues regulate cellular behavior, influence tissue development, and contribute to disease progression [78, 34, 3, 12, 25, 26, 46, 2, 48, 58]. Moreover, investigating how the microstructure, in terms of fiber and network properties, relates to the measured properties enables material behavior prediction and provides a solid foundation for designing engineered tissues with tailored macroscopic mechanics [11, 9].
In this study, we propose TopoGEN, a framework that integrates three-dimensional image-informed fiber network generation with non-linear finite element analysis to support the mechanistic investigation of structure-function relationships in soft matter. Our framework features a novel generative pipeline designed to create topologically rich and microstructurally diverse discrete fiber networks that replicate architectural variations induced by experimental conditions (for example, temperature or pH). As a case study, we examine how microstructural changes in simulated collagen gels influence macroscale mechanical behavior. We compare these simulations to experimentally studied reconstituted collagen networks, where similar topological variations are induced by polymerization temperature [43]. Unlike previous studies [88, 43, 24, 49, 21, 22, 75, 44], which varied individual parameters such as average connectivity at fixed or random fiber lengths, or concentration at fixed length and connectivity, TopoGEN enables simultaneous control of all these structural features. This comprehensive control supports a deeper understanding of how the complete microstructural organization affects the bulk mechanical response of soft biological materials. Specifically, to extend the analysis beyond experimentally accessible mesoscale parameters like concentration and connectivity, we leverage TopoGEN’s ability to generate microstructures with targeted variations, altering one feature while keeping others constant, to isolate the individual effects of key microstructural attributes, including fiber morphology and constituent mechanics. This capability is achieved through explicit control over critical parameters, including fiber connectivity, morphology, and concentration, during the generative process, as demonstrated in prior works [64, 53, 52, 24]. However, unlike these earlier approaches, TopoGEN imposes biologically motivated constraints on network architecture. It limits the number of connections per node to represent loosely connected structures where fibers interact only via cross-links or branches, excluding configurations with higher-order connectivity that are less representative of biological reality.
@@ 2.1 Fiber network generation
We model biological networks as structures of connected fibers. To generate these networks, we partition a three-dimensional cubic domain into randomly seeded Voronoi polyhedra and represent the fibers and cross-links as the tessellation and vertices, respectively. We then optimize this Voronoi-based fiber network using an iterative optimization technique to match key topological features, i.e., average connectivity and fiber length distribution, to those of experimentally measured in biological fiber networks.
Figure 1: Optimization workflow. Starting from an initial random configuration, the algorithm iteratively minimizes the system’s energy by removing fibers (move 1) or moving nodes (move 2) until the target distribution of connectivity and fiber length is achieved.
To transform our initial Voronoi-generated random fiber network into neighboring states, we introduce two types of moves illustrated in Figure 1: dilutive transformations, which randomly remove fibers from the network, and density-preserving transformations, which randomly move nodes within the cubic domain. We define cross-links as nodes with four-fold connectivity and branches as nodes with three-fold connectivity. Starting from the four-fold connectivity typical of Voronoi networks, dilution allows branch formation until the target connectivity is reached. The Poisson length distribution used to generate our initial network results in a number of short fibers, which are inconsistent with the experimentally measured target distribution. To address this, we assign higher removal probabilities to shorter edges during the dilutive transformations and iteratively adjust node positions to match the log-normal distribution during the density-preserving transformations. Once the target connectivity is achieved, we apply a simulated annealing procedure [96, 53, 52, 24] to optimize the fiber length distribution. We set the to-be-minimized annealing system energy as the distance of the initial microstructural configuration from the target microstructural configuration achieved by gradually cooling the network and allowing it to transition to new neighboring microstructures. These neighboring states are generated by applying small random perturbations to the configuration at each iteration until the current and target distributions match. To address the challenge of directly comparing continuous length distributions to the log-normal target [53], we implement a binning algorithm that partitions fiber lengths into disjoint intervals ($b$). Therefore, we assign each length observation ($x$) to a specific interval and compute a discretized cost function expressing the difference between the binned distribution and the target. To this aim, we implement the Kullback–Leibler (KL) divergence [47], also known as relative entropy ($\mathcal{D}_{KL}$), which is a non-symmetric measure quantifying how one probability distribution $p(x)$ diverges from a second reference probability $q(x)$ distribution:
@@
At each iteration, moves are accepted based on the Metropolis criterion [60], with a temperature-like parameter $T$ controlling the likelihood of accepting worse configurations. We set the temperature to decrease exponentially as $T=0.95^{k}\cdot T_{0}$, where $k$ is the iteration step [24, 64]. The initial temperature $T_{0}$ is set to accept  50% of moves that increase the energy, as in [24, 64]. Annealing stops when $T<1\times 10^{-4}\cdot T_{0}$, when the relative energy drop stays below $1\times 10^{-5}$ over 500 iterations, or when the maximum number of iterations is reached.
Our generative pipeline relies on four user-defined physical variables informed by the microstructure: fiber radius, target length distribution, average connectivity, and the volume of the RVE. The volume fraction, defined as the ratio of total fiber volume to total RVE volume, is a dependent variable influenced by these four parameters. To control the volume fraction, the user can adjust the number of seeds for Voronoi tessellation. To automate the selection of the optimal seed count that matches the target volume fraction, we propose a simple method. Since length optimization preserves density and only slightly affects volume fraction, it is reasonable to infer that connectivity optimization provides a reliable indication of how volume fraction responds to changes in the number of seeds. Starting from a user-provided seed estimate, a binary search algorithm determines the seed count that best approximates the desired volume fraction before launching the generative pipeline. Supplementary materials include additional analysis regarding the relationship between the fiber length and the number of seeding points (see Figure S.1).
@@ 2.2 Mechanical equilibrium at the mesoscale
Modeling soft tissues at the level of individual fibers (microscale) or the level of discrete fiber networks (mesoscale) is computationally impractical for large volumes undergoing complex loading and boundary conditions. Instead of extending our discrete fiber network model to the entire tissue scale (macroscale), we focus on a representative volume element (RVE) large enough to statistically capture the mechanical properties of the microstructure yet small enough to reduce the computational cost. Within this mesoscale domain, we define a constitutive behavior for each individual fiber in the network and apply macroscopic deformation gradients (i.e., tissue loading conditions) as boundary conditions. Solving this boundary value problem with the finite element method yields the macroscale’s local (i.e., for every finite element, at each integration point, within every time step, at each Newton iteration) constitutive law [32, 70]. In this study, we use the commercial finite element analysis software Abaqus [18] to retrieve the numerical solution of the RVEs boundary value problem. The transition from the RVE to the full tissue scale is only admissible if the Hill-Mandel condition [35] is satisfied, ensuring that the potential energy remains consistent when passing from the meso- to the macroscale. This condition is ideally met with infinitely large RVEs; however, with finite-sized RVEs, the consistency requires careful handling of the boundary conditions and the RVE size. This can be achieved through three approaches: (1) enforcing zero microfluctuations throughout the RVE, i.e., coupling local RVE deformation affinely to the macroscale deformation gradient; (2) setting zero microfluctuations only at the boundary via uniform displacement boundary conditions while allowing internal nodes to deform non-affinely (i.e., with deviations from a direct one-to-one correspondence to the macroscale deformation gradient); or (3) imposing periodicity of the microfluctuation field at the boundary via periodic boundary conditions (PBCs) and allowing internal nodes to deform non-affinely. In this work, we adopt PBCs since they are known to estimate the overall material properties better than the other alternatives mentioned [61, 87, 92, 45, 72] and lie closer to the effective properties compared to the upper and lower bounds defined by uniform displacement and traction boundary conditions. Another advantage of adopting PBCs is their ability to capture the effect of non-affinity on the RVE response by allowing internal nodes to move freely within our loosely connected networks. Given that biological networks are - in essence - not periodic structures, incorporating these PBCs is not a trivial task.
@@
Figure 2 highlights how we enforce periodicity at the boundary nodes while maintaining the internal microfluctuations. First, we randomly seed our RVE domain with N seeding points and generate 26 periodic replicas of this domain by offsetting the seeding points around the central domain (Figure 2, step 1). Next, we apply Voronoi tessellation using 27 X N points as seeding particles and extract the interconnected fibers from the skeletonized diagrams (Figure 2, step 2). We crop the larger network to the original cubic domain and assemble a periodic edge connectivity map between each boundary node and its periodic counterpart on the opposing edge (Figure 2, step 3). Finally, we implement PBCs by constraining the relative displacements of all node pairs, as defined in our assembled periodic boundary node connectivity table, to individual dummy nodes. These dummy nodes, unattached to any fiber in the model, serve as reference points used in the nodal equations that impose the periodicity of the nodally paired loading conditions [95]. In particular, at each individual dummy node, we enforce an affine macroscopic deformation gradient corresponding to the macroscopic RVE face deformation profile under study (e.g., uniaxial extension, biaxial extension, or shear, as illustrated in Figure 2, step 4). For each of these macroscopic deformation profiles, we introduce a linear interpolation map to constrain the boundary nodes to remain planar during shear or uni/biaxial tests.
Upon solving the mechanical equilibrium of our topologically optimized RVEs under various periodic loading conditions, we obtain the reaction forces on our PBC-constrained faces. We use these reaction forces to compute a representative stress tensor across the discrete fiber network by calculating the first Piola–Kirchhoff stress tensor ($\mathbf{P}$):
@@
where $A_{0}$ denotes the initial surface of our cubic RVE, ${N_{i}}$ are the dummy nodes, and $\mathbf{f}_{i}$ represents the resulting external forces acting on these nodes. We account for geometrical nonlinearity throughout the solution of these boundary value problems.
In our mechanical model, pipeline-generated node locations and their interconnections form the geometric input for finite element analysis. Here, one element represents a single fiber. Focusing purely on elastic effects, we assume fiber connections to be permanent, with no cross-link formation or branch breakage. We also consider that the persistence length of fibers, defined as the characteristic length along the polymer chain over which its direction remains correlated before bending, is typically larger than the contour length. As a result, the fibers are bending-stiff, and their thermal fluctuations can be neglected [10]. While the Euler-Bernoulli beam formulation could have been considered [84], neglecting shear stiffness is appropriate for fibers that are highly slender, with a cross-section-to-length ratio of less than 1:15. However, experiments in biological networks (e.g., collagen and fibrin) reveal a log-normal distribution of fiber lengths, which results in short fibers that violate this requirement [53, 52]. Therefore, we model individual fiber mechanics using second-order Timoshenko shear-flexible beams [69], which are particularly suited for representing biological networks [83]. We assume our fibers behave linearly elastic in both extension and compression, while accounting for fiber softening under compression. This asymmetry in the elastic modulus reflects the well-established observation that biological fibers soften their mechanical response under compression [66, 40]. Following the approach of [50], we implement this softening through a constitutive law applied to bending-resistant elements, while maintaining a linear elastic response in tension. As a result, our final model is bilinear, with the compressive elastic modulus set to one-tenth of the tensile modulus, consistent with [66]. To incorporate this bilinear tension-compression asymmetry, we developed a custom user-defined field USDFLD subroutine that defines the fiber elastic modulus at each integration point based on the axial strain of the loaded beam.